An electric iron is rated 500W, 200V calculate the :




(a) the current required to operate the iron




(b) the heat generated in 30minutes.





1
Expert's answer
2022-03-27T10:55:52-0400

(a)

"I=\frac{P}{V}=\frac{500\:\rm W}{200\:\rm V}=2.5\:\rm A"

(b)

"Q=Pt\\
=\rm 500\: W\times 30\times 60\: s=900\:000\:J"